### v1

@@ 1 Introduction
A core challenge in empirical aesthetics is linking subjective aesthetic responses (such as the experience of beauty or emotional arousal), to visual properties of the stimulus and their representations in the brain. In the case of visual aesthetics, a popular “bottom-up” approach assumes that responses are influenced by a combination of low-, mid-, and high-level image features, corresponding to stages in the visual hierarchy. Low-level features include contrast, lightness, saturation, color, etc. (Carandini2005, ); mid-level ones involve grouping, contours, segmentation, depth, etc. (Freeman2013, ; Roe2012, ; Welchman2005, ) and high-level features relate to object recognition and semantic associations (Goodale1992, ). These studies have shown that features such as color, contrast, saturation, aspect ratio, contour (Farzanfar2023, ; Iigaya2021, ; McManus1980, ; McManus1981, ; Reber1998, ), image quality, anisotropy, complexity and fractal self-similarity (Mallon2014, ; Spehar2003, ; Tinio2009, ) are modest predictors of aesthetic preference. Other features that seem to predict preferences are symmetry (Bertamini2019, ), which is preferred on abstract geometrical patterns (Jacobsen2002, ) and faces (Rhodes2006, ), and smoothly curved contours, which are preferred over angular contours (Bar2006, ; Bertamini2016, ; Clemente2023, ; Vartanian2013, ).
@@
Despite their success, the explanatory power of feature-based models is limited by small effect sizes, dataset constraints, strong inter-individual variability and methodological variability (for a review see Vessel et al (Vessel2022, )). Aesthetic judgments are shaped not only by perceptual input but also by context, memory, emotion and other cognitive factors that vary across individuals and cultures (Bignardi2021, ; Martinez2020, ; Palmer2010, ; Skov2022, ; Vessel2010, ). This entanglement of perceptual and cognitive factors complicates efforts to predict aesthetic responses from image features alone and is reflected in the strong confounding between shared versus individual visual appeal (Honekopp2006, ; Vessel2012, ; Vessel2013, ). In his book “A Brief Tour of Human Consciousness”, best-selling author V.S. Ramachandran states that “assuming that 90% of the variance seen in art is driven by cultural diversity, this is what people already study and it’s called Art History. As a scientist, I am interested in the 10% that is universal.” (Ramachandran2004, ). Many researchers that encountered this confounding found that observers do not often agree in what they find visually appealing. In a series of preference study of digital images, Vessel et al (Vessel2018, ; Vessel2010, ) found that observers’ judgments revealed a strong degree of shared taste for real-world scenes but were much more idiosyncratic in their ratings of novel, visually diverse images. In other words, the presence of higher-level, unshared semantic meanings and associations led to variability in ratings (Martinez2020, ). A similar result was found for color preferences, which were predicted by participant’s likings of objects associated to those colors (Palmer2010, ).
@@
In a rigorous examination of the sources of variance in aesthetic judgments—whether stemming from the stimuli, the observers, or their interaction—Martinez et al(Martinez2020, ) provided estimates that stabilized with repeated measures. Using variance component analysis (VCA) on a limited dataset comprising faces, single objects, and noise patterns, they found that observer-related factors accounted for most of the variability across all cases. However, stimulus characteristics contributed some variance in the case of faces, while stimulus-observer interactions contributed to the cases of both faces and objects. Importantly, they also noted that the results of such analyses depend heavily on the specificity of the hypotheses and the number of stimuli designed to target the relevant variance components.
@@ 2.1 Common methodological problems
The challenge of untangling perceptual and cognitive factors in the putative relationship between image features and aesthetic appreciation emerges already at the level of response measurement and the datasets used for image valuation. Response measurements can be broadly classified into direct and indirect types. Direct measures capture participants’ explicit judgments and preferences, such as binary responses (“yes”/“no”, “ugly”/“beautiful”), Likert scales, continuous rating scales, pairwise comparisons, and free-adjustment sliders (Belfi2019, ; Ishizu2011, ; Jacobsen2006, ; Kim2007, ; Vessel2019, ). Indirect measures reflect behavioral and physiological correlates of aesthetic engagement, including viewing time, reaction time, arousal levels, pupil dilation, heart rate, and economic proxies like “willingness to pay”, etc. (Aharon2001, ; Laeng2016, ; Lopez2020, ; Salimpoor2011, ; Smith2010, ; Tschacher2012, ). A further level of indirectness involves neurophysiological measurements, such as patterns of brain activity associated with different experimental conditions, typically captured via blood oxygen level–dependent (BOLD) signals in fMRI studies (Bar2007, ; Boynton1999, ; Iigaya2023, ; Jacobsen2006, ; Keefe2018, ; Pasupathy2002, ). Understandably, all the results were obtained from small collections of images gathered under each author’s criteria (often including art) and typically rated by a small group of participants. The results of direct measures are difficult to compare—not only because of the interplay between perceptual and cognitive factors, but also due to significant methodological variability. Indirect measures, on the other hand, aim to access unmediated aesthetic appeal—at least partially and without involving conscious reflection—but there is little consensus on which of these measures reliably indicate aesthetic response to visual stimuli. Another unresolved issue is whether incorporating direct measures (e.g., image ratings) during the collection of indirect data (e.g., brain imaging) introduces additional variability across studies. Interpretation is further complicated by the use of different questionnaires to assess aesthetic response—for instance, asking how beautiful, moving, attractive, or likeable an image is—each potentially tapping into distinct yet overlapping constructs. The use of artwork in such research further amplifies cognitive influences, as it is difficult to control for recognition of specific styles (e.g., Picasso, Miró). Such recognition can trigger semantic associations and bias aesthetic judgments, given that prior knowledge of an artwork or artist can strongly shape perceived aesthetic value (Kirk2009, ).
@@
The psychology and neuroscience methods discussed above often rely on simplified or artificial stimuli (e.g., disembodied faces, geometric shapes, patterns, etc.) and small, arbitrarily selected image sets, prioritizing experimental control at the expense of ecological validity. This limitation can be addressed through the use of computer vision techniques and large-scale datasets containing thousands of internet images (Chounchenani2025, ; Datta2006, ; Jiang2010, ; Joshi2011, ; Liu2017, ; Murray2012, ). However, identifying consistent correlations between image features and aesthetic judgments in such datasets remains extremely challenging. The same difficulty applies to training AI models to predict aesthetic responses. This reveals an apparent paradox: AI models require large datasets rated by many observers, yet as the datasets become broader and more general, the entanglement between perceptual and cognitive factors in observer responses increases—making it harder to disentangle and model the sources of aesthetic judgment. To address this, computer-based studies often classify the dataset content in a small number of categories (“ugly”/ “beautiful”), create overly-complex models, and sometimes consider an arbitrarily small fraction of the images contained in the datasets, without controlling for the probability of committing a type I error.
@@
Large image datasets (LIDs) are typically sourced from a limited number of internet platforms, with the most commonly used being Photo.net (photonet1993, ), Flickr (Flickr2003, ) and DPChallenge (DPChallenge2016, ). These platforms contain the tools necessary for users to upload digital images and cast votes on different aspects of the images uploaded by other users. Table 1 shows a list of some of the most successful image datasets used in computational aesthetics at the time of writing.
@@
Photo.net Photo database (Datta2006, ; Joshi2011, ) 14,839 Assessed by an average of 10 observers
Photo.net Photo database (Datta2010, ) 20,000 Assessed by at least 10 observers
Photo.net Waterloo-IAA (Liu2017, ) 1,100 Assessed for beauty in a lab-controlled user study (26 observers)
DPChallenge.com DPChallenge (Datta2006, ; Joshi2011, ) 16,509 Assessed by an average of 205 observers
DPChallenge.com AVA (Murray2012, ) 250,000 Assessed by an average of 200 observers
DPChallenge.com CUHK (Yan2006, ) 60,000 Assessed by at least 100 observers
DPChallenge.com PhotoQuality (CUHK-PQ) (Luo2011, ) 17,690 Assessed by an average of 200 observers and manually labelled using binary ratings
Flickr.com AROD (Schwarz2018, ) 320,000 Assessed through time-independent statistics (“faves” and “views”) by thousands of observers
Flickr.com FAE-Captions (Jin2022, ) 251,000 Assessed through image descriptions
Flickr.com FLICKR-AES (Ren2017, ) 40,000 Assessed by 210 unique Amazon Mechanical Turk (AMT) annotators
Other (Mix) IAD (Lu2015, ) 1,500,000 Collected from DPChallenge (300,000 images) and PHOTO.NET (1,200,000 images)
Other (Mix) Kodak Aesthetics (Jiang2010, ) 1,500 Collected from a number of sources (Flickr, Kodak Picture of the Day, study observers, etc.) and assessed by 30 observers
Other Terragalleria (Joshi2011, ) 14,449 Contains images of travel photography collected by Dr Quang-Tuan Luong and assessed by an average of 20 observers
@@
  • Evaluation methods are diverse and have been implemented using a range of paradigms: namely Absolute Category Rating (observers provide a score based on a fixed set of predefined categories), Likert Scale ratings (observers assess images on a scale of discrete values, usually ranging from 1 to 10) and Automatic Sentiment Analysis of the image comments and descriptions (see for example Jin et al(Jin2022, ); Ren et al(Ren2017, ) and Schwarz et al(Schwarz2018, )), which is common in databases derived from Flickr.com. Furthermore, platforms include ratings on several dimensions such as originality, technical merit, aesthetic content, or general appeal and the scores of each category are then averaged to derive an overall aesthetic rating (see for example Datta et al(Datta2006, )). These methodological differences may become an issue in studies that mix images from different sources (see for example Jiang et al(Jiang2010, ), Joshi et al(Joshi2011, ) and (Lu2015, )).
@@
  • Observer tasks are not uniform and may include biases (Joshi2011, ). Tasks consist in most cases of scoring, but also some kind of description, and in some cases the inclusion of the image in ’favorites’ or its sharing with others (e.g. Flickr.com). In many cases, it remains unclear whether participants are judging aesthetic beauty, technical quality, or both, as instructions are often vague or unspecified (Joshi2011, ). For example, what is considered “interesting” in one platform (Flickr.com) may not be treated as being “aesthetically pleasing” in another (Photo.net).
@@
  • The availability of highlights, descriptions, comments, and ratings prior to voting fosters “preference groups” influenced by prestigious or expert users. Experts, for instance, show a stronger correlation between originality and quality (Hekkert1996, ) and tend to value challenging, emotionally provocative works more highly (Winston1992, ).
@@
  • Some studies rely on binary distributions: Images are classified in only two categories (High/Low) to obtain high classification accuracy scores (see for example Datta et al., (Datta2006, ; Datta2010, ; Lu2015, ; Luo2011, ; Wang2019, ; Yan2006, ). This represents an acknowledgment of the limitations of the methods and restricts the understanding one can derive from these results.
@@
  • Some studies include arbitrarily chosen subsets, e.g. certain “challenges” (Iigaya2021, ) or high-consensus images (Luo2011, ; Yan2006, ). Others extract a large number of features (up to 1000) and calculate correlations without applying correction methods, such as the Bonferroni correction, to control for Type I errors (see for example Liu & Wang, (Liu2017, )).
@@
  • Some studies rely on a limited number of images or observers (see for example Bhattacharya et al(Bhattacharya2010, ), Iigaya et al(Iigaya2021, ), Jiang et al (Jiang2010, ), Liu & Wang, (Liu2017, )). Although there is no established consensus on the minimum required sample size, the high variability in responses—particularly in semantically rich datasets—indicates that robust analyses demand several thousand images and a substantial pool of observers to yield meaningful and generalizable results.
@@
The purpose of outlining these limitations is not to undermine previous studies’ contributions, but to underscore the inherent challenges in applying methodologies centered on low- and mid-level feature extraction to a task that is predominantly influenced by high-level, cultural, and context-dependent factors. These points are not new and were thoroughly analyzed in a seminal work by Joshi et al(Joshi2011, ) which also examined other shortcomings of popular platforms at the time (Photo.net, DPChallenge.com and Terragalleria), e.g. their idiosyncratic distributions of average aesthetics scores, the presence of ’challenges’, and their bias towards beauty, which are discussed below.
@@ 2.2 Problems derived from other inherent database biases
In most platforms, images are submitted under specific labels (“titles”, “themes” or “challenges”), a practice that sets a semantic background likely to affect subsequent aesthetic judgements. For instance, contributors to DPChallenge.com may evaluate an image by how close it fits to the “challenge” title, i.e., they might favor an image belonging to “Bad hair day” because it contains a pet with unruly hair or may prefer an image within “Through the eyes of a child” because it contains their favorite childhood toys. In other cases, the challenge’s winner decision may be decided on whether an image plays a humoristic twist on the issues of the day. In computational learning this is referred as “semantic gap” (Smeulders2000, ), meaning the discrepancy between the information extracted from the visual data and the interpretation that observers make of the same data in a given context. Despite sharing certain structural features, all these databases contain large assemblies of objects with various semantic interpretations, depending on the situation, challenge, title, theme, etc. This contextual content acts as a strong confounding factor when it comes to predicting aesthetic valuations from the image visual content.
@@
Another major drawback of these databases is the small variability in the aesthetic dimension of the images contained and their arguable bias towards beauty (Joshi2011, ). The low variability is illustrated by plotting the histogram of the number of images in each aesthetic valuation interval for the AVA dataset (Murray2012, ). Figure 1 shows that images rated as “very ugly” (aesthetic valuation in the 0-1 range) or “very beautiful” (aesthetic valuation in the 9-10 range) are statistically rare, with most ratings in the range [4.5-6.5]. This pattern may be attributed to high inter-observer variability: extreme ratings are often offset by more moderate ones, and according to the law of large numbers, such variations tend to cancel each other out. As a result, the average aesthetic score for most images is drawn toward the center of the scale. This leads to a dataset with a limited dynamic range in the aesthetic dimension. In other words, despite the diversity of image content, there are very few images that elicit widespread agreement as either "very ugly" or "very beautiful." Consequently, the overall distribution of scores approximates a normal curve with a relatively narrow standard deviation.
@@
The bias towards beauty combined with the limited dynamic range of the database, makes computational learning problematic. Indeed, any correlation analysis using a reduced range for one of the variables is likely to lead to spurious statistical outcomes, whether it overestimates, underestimates, or simply misses the correlation considering the full range of value for the variable. Another possible statistical trap when only considering a reduced range of values is to miss non-monotonic relationships between image metrics and aesthetic valuations. Some aesthetical valuations can, for example, depend on predictors such as complexity, following an inverted-U shape relationship (Berlyne1971, ). Such a dependence would spuriously be reduced as a monotonic relationship when restricting the range of values for the predictor.

### v2

@@ 1 Introduction
A core challenge in empirical aesthetics is linking subjective aesthetic responses (such as the experience of beauty or emotional arousal), to visual properties of the stimulus and their representations in the brain. In the case of visual aesthetics, a popular “bottom-up” approach assumes that responses are influenced by a combination of low-, mid-, and high-level image features, corresponding to stages in the visual hierarchy. Low-level features include contrast, lightness, saturation, color, etc. [1]; mid-level ones involve grouping, contours, segmentation, depth, etc. [2, 3, 4] and high-level features relate to object recognition and semantic associations [5]. These studies have shown that features such as color, contrast, saturation, aspect ratio, contour [6, 7, 8, 9, 10], image quality, anisotropy, complexity and fractal self-similarity [11, 12, 13] are modest predictors of aesthetic preference. Other features that seem to predict preferences are symmetry [14], which is preferred on abstract geometrical patterns [15] and faces [16], and smoothly curved contours, which are preferred over angular contours [17, 18, 19, 20].
@@
Despite their success, the explanatory power of feature-based models is limited by small effect sizes, dataset constraints, strong inter-individual variability and methodological variability (for a review see Vessel et al [21]). Aesthetic judgments are shaped not only by perceptual input but also by context, memory, emotion and other cognitive factors that vary across individuals and cultures [22, 23, 24, 25, 26]. This entanglement of perceptual and cognitive factors complicates efforts to predict aesthetic responses from image features alone and is reflected in the strong confounding between shared versus individual visual appeal [27, 28, 29]. In his book “A Brief Tour of Human Consciousness”, best-selling author V.S. Ramachandran states that “assuming that 90% of the variance seen in art is driven by cultural diversity, this is what people already study and it’s called Art History. As a scientist, I am interested in the 10% that is universal.” [30]. Many researchers that encountered this confounding found that observers do not often agree in what they find visually appealing. In a series of preference study of digital images, Vessel et al [31, 26] found that observers’ judgments revealed a strong degree of shared taste for real-world scenes but were much more idiosyncratic in their ratings of novel, visually diverse images. In other words, the presence of higher-level, unshared semantic meanings and associations led to variability in ratings [23]. A similar result was found for color preferences, which were predicted by participant’s likings of objects associated to those colors [24].
@@
In a rigorous examination of the sources of variance in aesthetic judgments—whether stemming from the stimuli, the observers, or their interaction—Martinez et al[23] provided estimates that stabilized with repeated measures. Using variance component analysis (VCA) on a limited dataset comprising faces, single objects, and noise patterns, they found that observer-related factors accounted for most of the variability across all cases. However, stimulus characteristics contributed some variance in the case of faces, while stimulus-observer interactions contributed to the cases of both faces and objects. Importantly, they also noted that the results of such analyses depend heavily on the specificity of the hypotheses and the number of stimuli designed to target the relevant variance components.
@@ 2.1 Common methodological problems
The challenge of untangling perceptual and cognitive factors in the putative relationship between image features and aesthetic appreciation emerges already at the level of response measurement and the datasets used for image valuation. Response measurements can be broadly classified into direct and indirect types. Direct measures capture participants’ explicit judgments and preferences, such as binary responses (“yes”/“no”, “ugly”/“beautiful”), Likert scales, continuous rating scales, pairwise comparisons, and free-adjustment sliders [32, 33, 34, 35, 36]. Indirect measures reflect behavioral and physiological correlates of aesthetic engagement, including viewing time, reaction time, arousal levels, pupil dilation, heart rate, and economic proxies like “willingness to pay”, etc. [37, 38, 39, 40, 41, 42]. A further level of indirectness involves neurophysiological measurements, such as patterns of brain activity associated with different experimental conditions, typically captured via blood oxygen level–dependent (BOLD) signals in fMRI studies [43, 44, 45, 34, 46, 47]. Understandably, all the results were obtained from small collections of images gathered under each author’s criteria (often including art) and typically rated by a small group of participants. The results of direct measures are difficult to compare—not only because of the interplay between perceptual and cognitive factors, but also due to significant methodological variability. Indirect measures, on the other hand, aim to access unmediated aesthetic appeal—at least partially and without involving conscious reflection—but there is little consensus on which of these measures reliably indicate aesthetic response to visual stimuli. Another unresolved issue is whether incorporating direct measures (e.g., image ratings) during the collection of indirect data (e.g., brain imaging) introduces additional variability across studies. Interpretation is further complicated by the use of different questionnaires to assess aesthetic response—for instance, asking how beautiful, moving, attractive, or likeable an image is—each potentially tapping into distinct yet overlapping constructs. The use of artwork in such research further amplifies cognitive influences, as it is difficult to control for recognition of specific styles (e.g., Picasso, Miró). Such recognition can trigger semantic associations and bias aesthetic judgments, given that prior knowledge of an artwork or artist can strongly shape perceived aesthetic value [48].
@@
The psychology and neuroscience methods discussed above often rely on simplified or artificial stimuli (e.g., disembodied faces, geometric shapes, patterns, etc.) and small, arbitrarily selected image sets, prioritizing experimental control at the expense of ecological validity. This limitation can be addressed through the use of computer vision techniques and large-scale datasets containing thousands of internet images [49, 50, 51, 52, 53, 54]. However, identifying consistent correlations between image features and aesthetic judgments in such datasets remains extremely challenging. The same difficulty applies to training AI models to predict aesthetic responses. This reveals an apparent paradox: AI models require large datasets rated by many observers, yet as the datasets become broader and more general, the entanglement between perceptual and cognitive factors in observer responses increases—making it harder to disentangle and model the sources of aesthetic judgment. To address this, computer-based studies often classify the dataset content in a small number of categories (“ugly”/ “beautiful”), create overly-complex models, and sometimes consider an arbitrarily small fraction of the images contained in the datasets, without controlling for the probability of committing a type I error.
@@
Large image datasets (LIDs) are typically sourced from a limited number of internet platforms, with the most commonly used being Photo.net [55], Flickr [56] and DPChallenge [57]. These platforms contain the tools necessary for users to upload digital images and cast votes on different aspects of the images uploaded by other users. Table 1 shows a list of some of the most successful image datasets used in computational aesthetics at the time of writing.
@@
Photo.net Photo database [50, 52] 14,839 Assessed by an average of 10 observers
Photo.net Photo database [58] 20,000 Assessed by at least 10 observers
Photo.net Waterloo-IAA [53] 1,100 Assessed for beauty in a lab-controlled user study (26 observers)
DPChallenge.com DPChallenge [50, 52] 16,509 Assessed by an average of 205 observers
DPChallenge.com AVA [54] 250,000 Assessed by an average of 200 observers
DPChallenge.com CUHK [59] 60,000 Assessed by at least 100 observers
DPChallenge.com PhotoQuality (CUHK-PQ) [60] 17,690 Assessed by an average of 200 observers and manually labelled using binary ratings
Flickr.com AROD [61] 320,000 Assessed through time-independent statistics (“faves” and “views”) by thousands of observers
Flickr.com FAE-Captions [62] 251,000 Assessed through image descriptions
Flickr.com FLICKR-AES [63] 40,000 Assessed by 210 unique Amazon Mechanical Turk (AMT) annotators
Other (Mix) IAD [64] 1,500,000 Collected from DPChallenge (300,000 images) and PHOTO.NET (1,200,000 images)
Other (Mix) Kodak Aesthetics [51] 1,500 Collected from a number of sources (Flickr, Kodak Picture of the Day, study observers, etc.) and assessed by 30 observers
Other Terragalleria [52] 14,449 Contains images of travel photography collected by Dr Quang-Tuan Luong and assessed by an average of 20 observers
@@
  • Evaluation methods are diverse and have been implemented using a range of paradigms: namely Absolute Category Rating (observers provide a score based on a fixed set of predefined categories), Likert Scale ratings (observers assess images on a scale of discrete values, usually ranging from 1 to 10) and Automatic Sentiment Analysis of the image comments and descriptions (see for example Jin et al[62]; Ren et al[63] and Schwarz et al[61]), which is common in databases derived from Flickr.com. Furthermore, platforms include ratings on several dimensions such as originality, technical merit, aesthetic content, or general appeal and the scores of each category are then averaged to derive an overall aesthetic rating (see for example Datta et al[50]). These methodological differences may become an issue in studies that mix images from different sources (see for example Jiang et al[51], Joshi et al[52] and [64]).
@@
  • Observer tasks are not uniform and may include biases [52]. Tasks consist in most cases of scoring, but also some kind of description, and in some cases the inclusion of the image in ’favorites’ or its sharing with others (e.g. Flickr.com). In many cases, it remains unclear whether participants are judging aesthetic beauty, technical quality, or both, as instructions are often vague or unspecified [52]. For example, what is considered “interesting” in one platform (Flickr.com) may not be treated as being “aesthetically pleasing” in another (Photo.net).
@@
  • The availability of highlights, descriptions, comments, and ratings prior to voting fosters “preference groups” influenced by prestigious or expert users. Experts, for instance, show a stronger correlation between originality and quality [65] and tend to value challenging, emotionally provocative works more highly [66].
@@
  • Some studies rely on binary distributions: Images are classified in only two categories (High/Low) to obtain high classification accuracy scores (see for example Datta et al., [50, 58, 64, 60, 67, 59]. This represents an acknowledgment of the limitations of the methods and restricts the understanding one can derive from these results.
@@
  • Some studies include arbitrarily chosen subsets, e.g. certain “challenges” [7] or high-consensus images [60, 59]. Others extract a large number of features (up to 1000) and calculate correlations without applying correction methods, such as the Bonferroni correction, to control for Type I errors (see for example Liu & Wang, [53]).
@@
  • Some studies rely on a limited number of images or observers (see for example Bhattacharya et al[68], Iigaya et al[7], Jiang et al [51], Liu & Wang, [53]). Although there is no established consensus on the minimum required sample size, the high variability in responses—particularly in semantically rich datasets—indicates that robust analyses demand several thousand images and a substantial pool of observers to yield meaningful and generalizable results.
@@
The purpose of outlining these limitations is not to undermine previous studies’ contributions, but to underscore the inherent challenges in applying methodologies centered on low- and mid-level feature extraction to a task that is predominantly influenced by high-level, cultural, and context-dependent factors. These points are not new and were thoroughly analyzed in a seminal work by Joshi et al[52] which also examined other shortcomings of popular platforms at the time (Photo.net, DPChallenge.com and Terragalleria), e.g. their idiosyncratic distributions of average aesthetics scores, the presence of ’challenges’, and their bias towards beauty, which are discussed below.
@@ 2.2 Problems derived from other inherent database biases
In most platforms, images are submitted under specific labels (“titles”, “themes” or “challenges”), a practice that sets a semantic background likely to affect subsequent aesthetic judgements. For instance, contributors to DPChallenge.com may evaluate an image by how close it fits to the “challenge” title, i.e., they might favor an image belonging to “Bad hair day” because it contains a pet with unruly hair or may prefer an image within “Through the eyes of a child” because it contains their favorite childhood toys. In other cases, the challenge’s winner decision may be decided on whether an image plays a humoristic twist on the issues of the day. In computational learning this is referred as “semantic gap” [69], meaning the discrepancy between the information extracted from the visual data and the interpretation that observers make of the same data in a given context. Despite sharing certain structural features, all these databases contain large assemblies of objects with various semantic interpretations, depending on the situation, challenge, title, theme, etc. This contextual content acts as a strong confounding factor when it comes to predicting aesthetic valuations from the image visual content.
@@
Another major drawback of these databases is the small variability in the aesthetic dimension of the images contained and their arguable bias towards beauty [52]. The low variability is illustrated by plotting the histogram of the number of images in each aesthetic valuation interval for the AVA dataset [54]. Figure 1 shows that images rated as “very ugly” (aesthetic valuation in the 0-1 range) or “very beautiful” (aesthetic valuation in the 9-10 range) are statistically rare, with most ratings in the range [4.5-6.5]. This pattern may be attributed to high inter-observer variability: extreme ratings are often offset by more moderate ones, and according to the law of large numbers, such variations tend to cancel each other out. As a result, the average aesthetic score for most images is drawn toward the center of the scale. This leads to a dataset with a limited dynamic range in the aesthetic dimension. In other words, despite the diversity of image content, there are very few images that elicit widespread agreement as either "very ugly" or "very beautiful." Consequently, the overall distribution of scores approximates a normal curve with a relatively narrow standard deviation.
@@
The bias towards beauty combined with the limited dynamic range of the database, makes computational learning problematic. Indeed, any correlation analysis using a reduced range for one of the variables is likely to lead to spurious statistical outcomes, whether it overestimates, underestimates, or simply misses the correlation considering the full range of value for the variable. Another possible statistical trap when only considering a reduced range of values is to miss non-monotonic relationships between image metrics and aesthetic valuations. Some aesthetical valuations can, for example, depend on predictors such as complexity, following an inverted-U shape relationship [70]. Such a dependence would spuriously be reduced as a monotonic relationship when restricting the range of values for the predictor.